OpenStudy (anonymous):

What is the 7th term of the geometric sequence where a1 = 1,024 and a4 = -16?

OpenStudy (anonymous):

a4=ar^4-1 -16=1024r^3 thus r=-1/4

OpenStudy (anonymous):

-__- i didn't get that when i tried it

OpenStudy (anonymous):

7 th term a7=a1r^7-1 a7=1024(-1/4)^6 a7=-0.25 answer

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

whats the geometric series formula?

OpenStudy (anonymous):

an=a1r^n-1

OpenStudy (anonymous):

umm actually the one where you'd need to find the sum of everything

OpenStudy (anonymous):

sum s=a1(r^n-1)/r-1

OpenStudy (anonymous):

ohh ahaha i got the things inside the parenthesis switched around ahaha thanks

OpenStudy (anonymous):

For future reference, -0.25 is not correct.

OpenStudy (lilymq):

Thank you @etessa7416

OpenStudy (lilymq):

Etessa7416 is right, I just took the test. The correct answer is 0.25, without the negative :)